Please help me.
Solve the equation.
-6b+6+4b=3b+11
thanks

Answer by Earlsdon(6294)

-6b + 6 + 4b = 3b + 11 Subtract 3b from both sides.
-6b + 6 + 4b - 3b = 11 Collect like-terms.
-5b + 6 = 11 Subtract 6 from both sides.
-5b = 5 Divide both sides by -5
b = -1
Check the solution by substituting -1 for b in the original equation.